If 1.43g Ar are added to 2.97 atm He in a 2.00 L cylinder at 27.0 °C, what is the total pressure of the resulting gaseous mixture?

P total = _____ atm
Answer #1

According to Daltons law,the total pressure exerted by a mixture of two or more non-reacting gases in a definite volume is equal to the sum of partial pressures of the constituent gases

Total pressure, Patm = PHe + PAr

From Ideal gas equation, PV =nRT

Pressure of Ar gas, P = nRT/V = (mass/molar mass)*(RT/V)

                      = (1.43/40)*(0.0821*300/2)

                      = 0.440 atm

total pressure of the resulting gaseous mixture = PHe + PAr = 2.97 + 0.440 = 3.410 atm